Question 7 If V-20 V and R : 10 kΩ the current equals 2 mA 0.2 A 200 μΑ 20 mA
Question 8 If V-5V and l-33 mA, the power dissipated by the resistor equals 165mW 16.5W 1.65mW 16.5mW
Answer #1

7)

we know KVL is

Voltage = current*resistance

or

V = I*R

I = V/R = 20/10000 = 2 *10^(-3) = 2 mA

option a) is correct

8) R = V/I = 5/33m = 5000/33 = 151.51 ohm

Power = R.(I^2) = 151.51*((33m)^2) = 0.165 Watt

= 165mW

option a) is correct.

m is known as milli and its value is 10^(-3) or 0.001